Smithen Company, a wholesale distributor, has been operating for only a few months. The company sells three products—sinks, mirrors, and vanities. Budgeted sales by product and in total for the coming month are shown below based on planned unit sales as follows:

Units Percentage
  Sinks 600 50%
  Mirrors 300 25%
  Vanities 300 25%
  Total 1,200 100%
Product
Sinks Mirrors Vanities Total
  Percentage of total sales 50% 30 20 100%
  Sales $240,790.00 100% $145,200 100% $98,010.00 100% $484,000.00 100%
  Variable expenses 72,237.00 30% 116,160 80% 53,905.50 55% 242,302.50 50%
  Contribution margin $168,553.00 70% $29,040 20% $44,104.50 455 $241,697.50 50%
  Contribution margin per unit 280.92 96.80 147.02
  Fixed expenses 209,000.00
  Operating income 32,697.50

Break-even point in unit sales :

fixed expenses/overall CM ratio = 209,000/0.5 = 418,000

Break-even point in unit sales:

total fixed expenses/weighted-average CM per unit = 209,000/201.41* = 1,037.66 units

*(280.92 x 0.5) + (96.80 x 0.25) + (147.02 x 0.25)

Assume that actual sales for the month total $582,030 (1,400 units), with the CM ratio and per unit amounts the same as budgeted. Actual fixed expenses are the same as budgeted, $209,000. Actual sales by product are as follows: sinks, $196,645 (490 units); mirrors, $271,040 (560 units); and vanities, $114,345 (350 units).

1. Prepare a contribution format income statement for the month based on actual sales data. (Round your percentage answers to the nearest whole number.)

SMITHEN COMPANY

Contribution Margin Income Statement
Product
Sinks Mirrors Vanities Total
Percentage of total % % % %
% % % %
% % % %
$0 0 % $0 0 % $0 0 % 0 0 %
Operating income (loss) $0

2. Compute the break-even point in sales dollars for the month, based on the actual data. (Round your intermediate calculations to the nearest whole percent. Round your final answer to the nearest whole dollar.)

3. Calculate the break-even point in unit sales for the month, based on the actual data. (Round your final answer to the nearest whole number.)
Answer #1

Please find below table useful to compute desired results: -

A B F G H I J K L 1 2 3 Percentage Mirrors Vanities Total 0.5 30 Units 600 300 300 1200 0.25 0.25 E Particulars Given Product

F G H sinks =B12 % of total Mirrors =B34/H34 =B14 % of total =D34/H34 Vanities =B16 % of total Total =F34/H34 =B9 =B36/H36 =D

End results would be as follows: -

A BC F G H I J K L Mirrors Vanities Total 30 20 300 E Particulars Given Product Sinks Percentage of total sales Sales 240,790

E F G H KL 31 32 Actual Figures 33 Mirrors sinks 490 % of total 35% % of total Vanities 40% 350 % of total Total 25% 1,400 56